OpenStudy (anonymous):

The length of a rectangle is 7 centimeters less than three times its width. Its area is 20 square centimeters. Find the dimensions of the rectangle

OpenStudy (anonymous):

3x - 7 = 20. 20 + 7 = 27 27 / 3 = 9 x = 9